# A Survey on Open Information Extraction from Rule-based Model to Large Language Model

### v6

@@ 5.1 Pre-neural Model Era
ORTE Task: Text $\rightarrow$ Relational Triplet
@@
In the beginning, Open IE systems were developed to create a universal model capable of extracting relation triplets through shallow features, such as Part-of-Speech (POS) that do not have lexical information, for instance, characterizing a verb based on its context. Normally, traditional machine learning models, such as Naive Bayes Rish et al. (2001) and Conditional Random Field Sutton et al. (2012), are used to train on shallow featuresYates et al. (2007); Wu and Weld (2010); Zhu et al. (2009). Using only lexical features will lead to problems of incoherent and uninformative relations. Therefore, lexical features and syntactic features are used to mitigate such problems Schmitz et al. (2012); Qiu and Zhang (2014); Mausam (2016). Later, rule-based models take advantage of hand-written patterns and rules to match relations Fader et al. (2011); Akbik and Löser (2012). To extract relations in a fine-grained way, clause-based models determine the set of clauses and identify clause types before extracting relations Del Corro and Gemulla (2013); Schmidek and Barbosa (2014); Angeli et al. (2015).
5.2 Neural Model Era: Open Relational Triplet Extraction
ORTE Task: Text $\rightarrow$ Relational Triplet
5.2.1 Labeling
@@
RnnOIE Stanovsky et al. (2018) is the first neural method, which formulates the OpenIE task as a sequence labeling problem. It uses a Bi-LSTM transducer to process input features, including word embeddings, part-of-speech tags, and indicated predicates. A Softmax classifier tags a BIO label for each token, after which triplets are constructed. Since one sentence usually contains more than one relation triplet, many approaches propose to avoid encoding and labeling the same input several times. OpenIE6 Kolluru et al. (2020a) adopts a novel Iterative Grid Labeling (IGL) architecture to capture dependencies among extractions without re-encoding. MacroIEBowen et al. (2021) reformulates the OpenIE as finding maximal cliques from the graph. DetIE Vasilkovsky et al. (2022) casts the task as a direct set prediction problem. SMiLe-OIE Dong et al. (2022) improves the model in an information-source view, using GCNs and multi-view learning to incorporate constituency and dependency information and aggregating semantic features and syntactic features by concatenating BERT embedding and graph embeddings.
@@
The sequence labeling paradigm is characterized by its efficiency, making it computationally advantageous for large-scale text processing. It yields readily interpretable output, as each token associates itself with a specific role, such as subject, relation, or object. A notable limitation of this approach is its lack of a holistic view, as it treats tokens in isolation, potentially failing to capture global context and complex relationships that extend beyond single tokens. Additionally, its output format may not adequately represent the nuanced variability of natural language.
5.2.2 Sequence to Sequence Generation
@@
Cui et al. (2018) casts OpenIE as a sequence-to-sequence generation problem and proposes NeuralOIE, which is an encoder-decoder model generating tuples with placeholders as a sequence according to the given input sentence. To enlarge the vocabulary and reduce the proportion of generated unknown tokens, NeuralOIE uses the attention-based coping mechanism. LogicianSun et al. (2018), based on attention-based sequence-to-sequence learning, transforms input sentences into structured facts. It employs a restricted copy mechanism to decide between copying information from input or selecting predefined keywords, enhancing reliability. IMoJIEKolluru et al. (2020b) is a generative OpenIE model that produces variable and diverse extractions for a sentence. It uses an iterative memory, implemented with a BERT encoder, to keep track of previous extractions. The model employs an LSTM decoder to generate extractions one word at a time until an "EndOfExtractions" token is reached.
The sequence generation paradigm focuses on generating sequences of text to represent extracted information, affording a more expressive and context-rich output. It excels in capturing global context and complex relationships, as it considers the broader contextual information in the text. This approach is adaptable to various languages and domains, rendering it versatile in a range of applications. Nevertheless, it entails complexity in training, potentially demanding larger datasets and longer training times. The generated sequences may also exhibit noise or ambiguities, necessitating post-processing for refinement. The output format from sequence generation models can vary, which poses challenges for downstream applications requiring standardized output structures.
5.3 Neural Model Era: Two-Stage Open Relation Extraction
ORSE Task: Entities + Text $\rightarrow$ Relation Span
@@ 5.3 Neural Model Era: Two-Stage Open Relation Extraction
Taking advantage of the remarkable representation capability of PLMs such as BERT, many researchers refine the model architecture into two stages to achieve more effective extractions. $\textrm{Multi}^{2}\textrm{OIE}$ Ro et al. (2020) is a two-stage labeling method. Its first stage is to label all predicates upon BERT-embedded hidden states instead of locating predicates with syntactic features. The second stage is to extract the arguments associated with each identified predicate by using a multi-head attention mechanism. Gen2OIE Kolluru et al. (2022) extends to a generative paradigm operating in two stages, generating all possible relational predicates and relation triplet sequentially. QuORE Yang et al. (2022) is a framework to extract relations and detect non-existent relationships based on the argument-context queries generated in the first stage.
@@
Some researchers in this modularized setting are exploring various intermediate representations to enhance the performance of the pipeline. OIE@OIA Wang et al. (2022e) is an adaptable OpenIE system that employs the methodology of Open Information eXpression (OIX) by parsing sentences into Open Information Annotation (OIA) Graphs. It consists of two components: an OIA generator that converts sentences into OIA graphs and a set of adaptors, each designed for a specific OpenIE task, allowing for efficient and versatile information extraction. By using different intermediate representations, Chunk-OIE Dong et al. (2023) introduces the Chunk sequence (SaC) as an intermediate representation layer while Yu et al. (2022) introduce directed acyclic graph (DAG) as a minimalist intermediate expression.
5.4 Neural Model Era: Open Relation Clustering
[…]
Representation by Defined Features.
Early clustering methods represent relation instances in feature space with the help of explicit features from various types of information. Ru et al. (2017) compares the contributions of different sequential patterns, syntactic information and the combination of the above to the representation of relation instances, which are clustered by a hierarchical clustering algorithm Zhao et al. (2005). The result shows that sequential patterns and syntactic information are both beneficial to relation representation. With more fine-grind features, Lechevrel et al. (2017) select core dependency phrases to capture the semantics of the relations between entities. Since different relations in one sentence can be viewed as noises when clustering a certain relation, Elsahar et al. (2017) propose a more resilient approach based on the shortest dependency path instead of directly cutting irrelevant information in sentences.
With the development of pre-trained language models, contextualized semantics can be better represented. Before the clustering step, recent clustering-based approaches tend to optimize relation representations with different supervision signals instead of manually extracting features based on different rules.
@@ 5.4 Neural Model Era: Open Relation Clustering
Semantic Representation by PLM–unsupervised learning. Unlike the above OpenIE systems that follow Banko et al. (2007) to use unsupervised learning methods, RSN Wu et al. (2019) exploits existing labeled data and relational facts in knowledge bases during training. To narrow the representation gap between pre-defined relations and novel relations, RSN learns a relational similarity matrix that can transfer relation knowledge from supervised data to unsupervised data. Wang et al. (2022c) and Genest et al. (2022) introduce a similar unsupervised prompt-based algorithm, MatchPrompt, which clusters sentences by leveraging representations from masked relation tokens within a prompt template. Its superb performance against traditional unsupervised methods indicates that fully leveraging the semantic expressive power of pre-trained models is very important. IRLCWang et al. (2022d) presents a robust solution for relation representation that employs data augmentation to create additional examples.
Semantic Representation by PLM–semi-supervised learning. Without taking advantage of labeled datasets, SelfORE Hu et al. (2020) proposes a self-supervised learning method for learning better feature representations for clustering. SelfORE is composed of three sections: (1) encode relation instances by leveraging BERT Devlin et al. (2019) to obtain relation representations; (2) apply adaptive clustering based on updated relation representations from (1) to assign each instance to a cluster with high confidence. In this way, pseudo labels are generated. (3) pseudo labels from (2) are used as supervision signals to train the relation classifier and update the encoder in (1). Repeat (2). As mentioned above high-dimensional vectors need to be clustered in a more complex way; Zhao et al. (2021) argue that such high-dimensional vectors contain too much irrelevant information for relation clustering, such as complex linguistic information. They propose a relation-oriented model based on SelfORE with a similar unsupervised training part and a modified supervised part. Duan et al. (2022) also make use of the distance between labeled data with every cluster center as soft pseudo labels. Zhao et al. (2023) have proposed a method of active learning, which allows the model to identify and present difficult data for manual annotation during the clustering process. This method not only reduces the amount of data that needs to be labeled but also achieves better experimental results.
Hierarchical Information. Apart from labeled data, knowledge bases also benefit OpenIE by generating positive and negative instances. Datasets generated from distant supervision bring in spurious correlations Roth et al. (2013); Jiang et al. (2018); Chen et al. (2021a). Fangchao et al. (2021) conduct interventions derived by KB to entities and context separately to avoid spurious correlations of them to relation types. OHRE Zhang et al. (2021a) proposes a top-down hierarchy expansion algorithm to cluster and label relation instances based on the distance between the KB hierarchical structure. In this way, clusters of existing relations are labeled clearly, and novel relations can be labeled as children relations of existing relation labels.
@@ 5.5 Large Language Models Era
ORTE Task: Text $\rightarrow$ Relational Triplet
@@
Fine-Tuning and Few-Shot. Consequently, efforts have been made to fine-tune pre-trained LLMs or employ in-context learning prompting strategies to utilize and enhance the instruction-following ability of LLMs. For example, Lu et al. (2023) addresses open-world information extraction, including unrestricted entity and relation detection, as an instruction-following generative task, and develops PIVOINE, a fine-tuned information extraction LLM that generates comprehensive entity profiles in JSON format. To minimize the need for extensive fine-tuning of LLMs, Ling et al. (2023) proposes various in-context learning strategies for performing relation triplet generation to improve the instruction-following ability of LLMs, and introduces an uncertainty quantification module to increase the confidence in the generated answers. Qi et al. (2023a) proposes an approach of constructing a consistent reasoning environment by mitigating the distributional discrepancy between test samples and LLMs. This strategy aims to improve the few-shot reasoning capability of LLMs on specific OpenIE tasks.
UIE and Applications. Besides reviewing the work that utilizes LLMs to address OpenIE, we then broaden our scope to 1. introduce some emerging trends and paradigms in universal information extraction (UIE), 2. exploration of how LLMs are applied to general IE tasks, and 3. the integration of LLMs in IE system pipelines in Appendix C.
We believe that this broader perspective provides readers with a comprehensive understanding of current trends and future directions in OpenIE and generic IE in the LLM era, enhancing their understanding of the field’s evolving dynamics, with the impact further discussed in Section 6.2.
@@ 6 Discussion
6.1 Source of information
In this section, we provide an overview of the information sources utilized by OpenIE models.
@@ 6.1 Source of information
Input-based information refers to features explicitly or implicitly present in the input unstructured text. Early OpenIE models extensively utilized explicit information such as shallow syntactic information, including part of speech (POS) tags and noun-phrase (NP) chunks Banko et al. (2007); Wu and Weld (2010); Fader et al. (2011). Although this approach is reliable, it does not capture all relation types Stanovsky et al. (2018), leading to the increasing use of deep dependency information, which reveals word dependencies within sentences Vo and Bagheri (2018); Elsahar et al. (2017). Subsequent OpenIE models have emphasized the use of semantic information to grasp literal meanings and linguistic structures, thereby enhancing the expression of relations despite the risk of over-specificity Vashishth et al. (2018); Wu et al. (2018). Recent models, including pre-trained language models, combine syntactic and semantic information to improve accuracy Hwang and Lee (2020); Ni et al. (2021). Further details are available in Appendix F.1.
@@
External information supplements OpenIE systems to enhance model performance. Early systems employ expert rules, including heuristic rules that integrate domain knowledge and assist in error tracing and resolution, based on syntactic analyses such as POS-tagging Chiticariu et al. (2013); Fader et al. (2011). Following this, the integration of hierarchical information from knowledge bases (KBs) has advanced knowledge representation learning. This integration provides structured hierarchies and detailed factual knowledge, which support more organized relation extraction and data augmentation Xie et al. (2016); Zhang et al. (2021a); Fangchao et al. (2021). More recently, with the development of LLMs, the pre-trained knowledge within these models is utilized, encapsulating extensive relational data Jiang et al. (2020); Petroni et al. (2020) and enabling efficient retrieval with well-designed instructions. Further details in Appendix F.2.

### v7

@@ 5.1 Pre-neural Model Era
In the beginning, OpenIE systems were developed to create a universal model capable of extracting relation triplets through shallow features, such as Part-of-Speech (POS) that do not have lexical information, for instance, characterizing a verb based on its context. Traditional machine learning models, such as Naive Bayes Rish et al. (2001) and Conditional Random Field Sutton et al. (2012), are used to train on shallow features Yates et al. (2007); Wu and Weld (2010); Zhu et al. (2009). Using only lexical features will lead to problems of incoherent and uninformative relations. Therefore, lexical features and syntactic features are used to mitigate such problems Schmitz et al. (2012); Qiu and Zhang (2014); Mausam (2016). Later, rule-based models take advantage of hand-written patterns and rules to match relations Fader et al. (2011); Akbik and Löser (2012). To extract relations in a fine-grained way, clause-based models determine the set of clauses and identify clause types before extracting relations Del Corro and Gemulla (2013); Schmidek and Barbosa (2014); Angeli et al. (2015).
5.2 Neural Model Era
@@
Sequence Labeling. RnnOIE Stanovsky et al. (2018) is the first neural method, which formulates ORTE task as a sequence labeling problem where inputs a sequence of tokens $\{x_{1},x_{2},...,x_{n}\}$ and outputs a sequence of labels $\{l_{1},l_{2},...,l_{n}\}$ with the same length $n$ as input. RnnOIE uses a Bi-LSTM to process input features, including word embeddings, POS tags, and indicated predicates. A Softmax classifier tags a BIO label for the last layer hidden state of each token, after which relation triplets are constructed. Since one sentence usually contains more than one relation triplet, many approaches propose to avoid encoding and labeling the same input several times Kolluru et al. (2020a); Bowen et al. (2021); Vasilkovsky et al. (2022). SMiLe-OIE Dong et al. (2022) improves the model in an information-source view, using GCNs and multi-view learning to incorporate constituency and dependency information and aggregating semantic features and syntactic features by concatenating BERT and graph embeddings. SMiLe-OIE aggregates semantic and syntactic features, which many other OpenIE systems omit.
@@
The sequence labeling paradigm is characterized by its computational efficiency, especially for large-scale text processing. It yields readily interpretable output, as each token associates itself with a specific role, such as subject, relation, object, spatial information, etc. It is limited by treating tokens in isolation, potentially failing to capture global context and complex relationships that extend beyond single tokens or cross sentences. Additionally, its output format may not adequately represent the nuanced variability of natural language.
@@
Sequence to Sequence Generation. Cui et al. (2018) casts OpenIE as a sequence-to-sequence (S2S) generation problem and proposes NeuralOIE, an encoder-decoder model generating a sequence of relation triplets conditioned by the input sentence. Facing unknown token openness problem, NeuralOIE uses the attention-based coping mechanism to enlarge the vocabulary. IMoJIEKolluru et al. (2020b) is an iterative generative OpenIE model that uses a BERT encoder to keep encoding previous generated relation triplets and generates the next triplet with an LSTM decoder until an "EndOfExtractions" token is reached. Compared with other neural models, IMoJIE generates a variable number of diverse extractions.
The S2S paradigm excels in capturing complex relationships, as it considers the broader contextual information. It is adaptable to various languages and domains but more demand in datasets and training times. The flexible output, while better-addressing openness challenges, also poses challenges for downstream applications requiring standardized output structures.
@@
Two-Stage Open Relation Extraction. Taking advantage of the remarkable representation capability of PLMs such as BERT, many researchers refine the model architecture into two stages to achieve more effective extractions. $\textrm{Multi}^{2}\textrm{OIE}$ Ro et al. (2020) is a two-stage labeling method. Its first stage is to label all predicates upon BERT-embedded hidden states instead of locating predicates with syntactic features. The second stage is to extract the arguments associated with each identified predicate by using a multi-head attention mechanism. In contrast to OpenIE systems that are limited to English, $\textrm{Multi}^{2}\textrm{OIE}$ is distinguished by its capability to process multilingual text. The intermediate representation can be other formats such as chunk sequence Kolluru et al. (2022) and directed acyclic Yang et al. (2022).
@@
Various intermediate representations are used to enhance the pipeline’s performance. OIE@OIA Wang et al. (2022d) is an adaptable OpenIE system that employs the method of Open Information expression (OIX) by parsing sentences into Open Information Annotation (OIA) Graphs. It consists of two components: an OIA generator that converts sentences into OIA graphs and a set of adaptors that trained to for versatile extraction formats. By using different intermediate representations, Chunk-OIE Dong et al. (2023) introduces the Chunk sequence (SaC) as an intermediate representation layer while Yu et al. (2022) introduces directed acyclic graph (DAG) as a minimalist intermediate expression.
@@
Open Relation Clustering. The clustering-based approaches are divided into relation representation and clustering. Some studies label clusters: Wang et al. (2022c) and Genest et al. (2022) introduce an unsupervised prompt-based algorithm, MatchPrompt, which clusters sentences by leveraging representations from masked relation tokens within a prompt template. Its superb performance against traditional unsupervised methods indicates that leveraging the semantic expressive power of pre-trained models is very important.
SelfORE Hu et al. (2020) propose a self-supervised learning method for learning better feature representations for clustering. SelfORE is composed of three sections: (1) encode relation instances by leveraging BERT Devlin et al. (2019) to obtain relation representations; (2) apply adaptive clustering based on updated relation representations from (1) to assign each instance to a cluster with high confidence. In this way, pseudo labels are generated. (3) pseudo labels from (2) are used as supervision signals to train the relation classifier and update the encoder in (1). Repeat (2) until converge. Compared with other OpenIE systems, it reduces dependence on labeled data. Based on similar self-supervised approaches, many works propose to reduce irrelevant information in relation representation Zhao et al. (2021), create pseudo labels Duan et al. (2022), and introduce human intervention during training to address the challenge of poorly clustered samples Zhao et al. (2023). During relation clustering, using complete input (sentence) representations as relation representations often leads to a significant decline in clustering performance when multiple relations exist within a single input. Semi-supervised learning has shown the best results, but the effectiveness largely depends on the quality of the annotated data.
Apart from labeled data, knowledge bases also benefit OpenIE by generating positive and negative instances. OHRE Zhang et al. (2021b) proposes a top-down hierarchy expansion algorithm to cluster and label relation instances based on the distance between the KB hierarchical structure. Existing relations are labeled with KB elements, and novel relations are labeled as children relations of existing ones. Using a structured KB can determine the broad category of a cluster’s relations, partially addressing cluster labeling issues. The KB structure can also define relation boundaries during clustering. However, errors in the KB can affect clustering accuracy, and building a high-quality KB still requires significant human effort.
A comparison of methods in ORTE, ORSE, and ORC settings reveals the strengths of task-specific strategies in achieving top performance. In ORTE, MacroIE excels with the highest F1 (54.8) on CaRB, leveraging maximal clique discovery, while SpanOIE leads on OIE16 and Re-OIE16 with F1 scores of 69.4 and 77.0. In ORSE, DragonIE tops CaRB with an F1 of 55.1, due to its efficient formulation, while OIE@OIA dominates on OIE16 and Re-OIE16 by using multi-task learning. For ORC, CaPL outperforms others on FewRel and TACRED through high-quality pseudo-labels, with MatchPrompt as a strong second, driven by its prompt-based framework.
@@ 5.3 Large Language Models Era
Fine-Tuning and Few-Shot. Consequently, efforts have been made to fine-tune pre-trained LLMs or employ in-context learning prompting strategies to utilize and enhance the instruction-following ability of LLMs. For example, Lu et al. (2023) addresses open-world information extraction, including unrestricted entity and relation detection, as an instruction-following generative task, and develops PIVOINE, refined via the process of instruction tuning, demonstrates the capacity to accurately extract pertinent entity information in accordance with directives articulated in natural language. Such a capability endows it with the flexibility to adapt to various information extraction tasks and requirements. To minimize the need for extensive fine-tuning of LLMs, Ling et al. (2023) proposes various in-context learning strategies for performing relation triplet generation to improve the instruction-following ability of LLMs, and introduces an uncertainty quantification module to increase the confidence in the generated answers. Qi et al. (2023a) proposes to construct a consistent reasoning environment by mitigating the distributional discrepancy between test samples and LLMs. This strategy aims to improve the few-shot reasoning capability of LLMs on specific OpenIE tasks.
@@ 6 Discussion
6.1 Co-Evolution of OpenIE: Task Settings and Model Capabilities
In this section, we unveil the connection between task settings and model capabilities in handling various features and information, demonstrating the intertwined development of both aspects.
@@ 6.1 Co-Evolution of OpenIE: Task Settings and Model Capabilities
Input-based information refers to features explicitly or implicitly present in the input unstructured text. Early OpenIE models extensively utilized explicit information such as shallow syntactic information, including part of speech (POS) tags and noun-phrase (NP) chunks Banko et al. (2007); Wu and Weld (2010); Fader et al. (2011). This approach is reliable, yet it does not capture all relation types Stanovsky et al. (2018), leading to the increasing use of deep dependency information, which reveals word dependencies within sentences Vo and Bagheri (2018); Elsahar et al. (2017). Subsequent OpenIE models have emphasized the use of semantic information to grasp literal meanings and linguistic structures, thereby enhancing the expression of relations despite the risk of over-specificity Vashishth et al. (2018); Wu et al. (2018). Recent models, including pre-trained language models, combine syntactic and semantic information to improve accuracy Hwang and Lee (2020); Ni et al. (2021). Further details in Appendix E.1.
@@
External information supplements OpenIE systems to enhance model performance. Early systems employ expert rules, including heuristic rules that integrate domain knowledge and assist in error tracing and resolution, based on syntactic analyses like POS-tagging Chiticariu et al. (2013); Fader et al. (2011). Following this, the integration of hierarchical information from knowledge bases (KBs) advances knowledge representation learning. This integration provides structured hierarchies and detailed factual knowledge, supporting more organized relation extraction and data augmentation Xie et al. (2016); Zhang et al. (2021b); Fangchao et al. (2021). With the developments of LLMs recently, the pre-trained knowledge within these models is utilized, encapsulating extensive relational data Jiang et al. (2020); Petroni et al. (2020) and enabling efficient retrieval with well-designed instructions. The strong generalization capabilities of LLM-based approaches allow them to embrace open-world knowledge, making them more robust and adaptable to various tasks and real-world applications. Further details in Appendix E.2.
6.2 Transforming OpenIE: The Impact of LLMs
When comparing the performance of LLMs with pre-LLM approaches, we see that LLMs have significantly advanced the task of OpenIE, often outperforming traditional methods. Zero-shot LLMs have achieved impressive and state-of-the-art (SOTA) results in various scenarios when evaluated on classic metrics such as token-level scorers Li et al. (2023a); Wang et al. (2022b). However, these models struggle with long-tail and more challenging tasks Gao et al. (2023). A major challenge for LLMs, compared to pre-LLM approaches like sequence tagging, is the issue of hallucination, which frequently occurs in various natural language generation tasks Ji et al. (2023), making faithfulness and reliability significant concerns. Traditional generative-based approaches from the pre-LLM era also suffer from errors such as redundant and incorrect extractions Schneider et al. (2017); Zhou et al. (2022), known as intrinsic hallucination. In contrast, LLM-based methods face the risk of both intrinsic hallucination and generating information unsupported by the original context or additional references, known as extrinsic hallucination Zhu et al. (2023); Ren et al. (2023); Li et al. (2023a). Despite these challenges, few-shot learning and fine-tuning can help mitigate issues related to long-tail challenges and hallucination to some extent. Additionally, until fundamental improvements in LLMs fully address these shortcomings, incorporating traditional approaches as supplementary supervisors when using LLM-based methods could potentially enhance reliability.
We also observe trends in developing universal paradigms for tackling various IE tasks. Recent advancements and the robust generalization capabilities of LLMs have led to the exploration of universal frameworks designed to address all IE tasks (UIE). These frameworks aim to leverage the shared capabilities inherent in IE, while also uncovering and learning from the dependencies between various tasks Xu et al. (2023b). This approach marks a significant shift from focusing on isolated subtasks, such as OpenIE, to a more integrated methodology that seeks a comprehensive understanding of the domain. The prevailing trajectory in developing universal IE frameworks is to establish unified, structured schemas, either natural language-based Wang et al. (2022a); Lu et al. (2022); Lou et al. (2023) or code-based Li et al. (2023d); Guo et al. (2023b); Sainz et al. (2023), to transform various IE tasks into a uniform task of structural information extraction while preserving the flexibility to adapt to the unique aspects of different tasks. More details on these approaches are provided in Appendix B.
Is OpenIE research going to its ends? LLMs bridge the gap between standard IE and OpenIE. LLMs are naturally suited for OpenIE, even under zero-shot scenarios, as they address both standard IE and OpenIE within the same task setting. In this setting, schemas and templates are designed to extract desired structural information. The primary difference is that standard IE schemas include more restrictions to limit the set of relations and entities. The flexibility and strong performance of LLMs in tackling various IE tasks through zero-shot and few-shot prompting, without requiring model updates, is attributed to their robust generalization ability acquired through pre-training. With this generalization capability, addressing both standard IE and OpenIE may not require fundamentally different methods; the main distinction lies in schema design. This significantly blurs the boundaries between standard IE and OpenIE. In the future, OpenIE might be viewed as a more complex and challenging scenario within IE tasks, rather than being distinctly separate from standard IE. Though we refrain from making a definitive conclusion, we can foresee OpenIE potentially merging into the broader scope of standard IE.